given the position function s=-t3+6t2 , t equal or more than 0


find the maximum velocity

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

v=ds/dt =-3t^2 +12t=0,t(12-3t), t=0 or 4, d2s/dt2 =-3t+12-3t=12-6t, second derivate is negative if t=4, therefore maximum velocity= 6t^2-t^2 occurs when t=4